OpenStudy (anonymous):

Find the intervals of increase or decrease: C(x)=x^(1/3)(x+4)

OpenStudy (anonymous):

\[C'(x)=\frac{ 1 }{ 3 } x^{-2/3}(x+4)+x ^{1/3}x\] \[C'(x)=x^\frac{ -2 }{ 3 }[\frac{ 1 }{ 3 }(x+4)+2x]\] \[=\frac{ (x+4)+2x }{ 3\sqrt[3]{x^2} }\] \[=\frac{ 3x+4 }{ 3\sqrt[3]{x^2} }\] x=-4/3

OpenStudy (anonymous):

This doesn't seem right though and I'm pretty sure it has something to do with satan's fractional exponents. Can someone show me where I went wrong?

OpenStudy (anonymous):

derivative of x+4 is not x

OpenStudy (anonymous):

your first line

OpenStudy (anonymous):